OpenStudy (anonymous):

What is the slope of the line x = –5?

OpenStudy (anonymous):

THis is a vertical line. Vertical lines have what kind of slope?

OpenStudy (anonymous):

Calcmathlete, i dont know :(

OpenStudy (anonymous):

 Alright. Let's test it out.

OpenStudy (anonymous):

These are the choices undefined 0 –5 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[m = \frac{y_1 - y_2}{x_1 - x_2}\]Two points on this line are (-5, 0) and (-5, 2)

OpenStudy (anonymous):

\[m = \frac{0 - 2}{-5 - (-5)} = \frac{-2}{0}\]The denominator can never be 0. Therefore it is undefined :)

OpenStudy (anonymous):

I thought so. Thank you for the explaination!

OpenStudy (anonymous):

np :) You might also want to keep in mind that horizontal lines have a slope of 0.

OpenStudy (anonymous):

Okay, i willl remember that ! (: